This is from a General Factoring Section...
Find Solutions?
-17w=8w^2+5w-40

Answer by scott8148(6628)

adding 17w __ 0=8w^2+22w-40

dividing by 2 __ 0=4w^2+11w-20

factoring __ 0=(4w-5)(w+4)

0=4w-5 __ 5/4=w

0=w+4 __ -4=w